A thin-walled hollow cylinder is rolling on a surface. What fraction of its total kinetic energy is in the form of rotational kinetic energy about the center of mass?

The answer is not 2/5.
Answer #1

thin-walled hollow cylinder means moment of inertia , I is given as

I = mr2

v = rw ( no slip condition)

so,

rotational kinetic energy, KR = 1/2 * I * w2

KR = 1/2 * mr2 * v2 / r2

KR = 1/2mv2

so,

fraction = KR / KT where KT is total kinetic energy

fraction = 1/2mv2 / 1/2mv2 + 1/2mv2

fraction = 1/2